\begin{thebibliography}{44}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{noa(2023)}]{noauthor_google_2023}
 2023.
\newblock \href {https://sscnr.net.in/google-bard-ai/} {Google {Bard} {AI},
  login, {Launch} date, {How} {To} {Use}, {Feature} {Full} {Detail} » sscnr}.
\newblock Section: Technology.

\bibitem[{Ali et~al.(2023)Ali, Tang, Connolly, Fridley, Shin, Sullivan, Cielo,
  Oyelese, Doberstein, Telfeian, Gokaslan, and Asaad}]{ali_performance_2023}
Rohaid Ali, Oliver~Y. Tang, Ian~D. Connolly, Jared~S. Fridley, John~H. Shin,
  Patricia L.~Zadnik Sullivan, Deus Cielo, Adetokunbo~A. Oyelese, Curtis~E.
  Doberstein, Albert~E. Telfeian, Ziya~L. Gokaslan, and Wael~F. Asaad. 2023.
\newblock \href {https://doi.org/10.1101/2023.04.06.23288265} {Performance of
  {ChatGPT}, {GPT}-4, and {Google} {Bard} on a {Neurosurgery} {Oral} {Boards}
  {Preparation} {Question} {Bank}}.
\newblock Pages: 2023.04.06.23288265.

\bibitem[{Anthropic()}]{anthropic_introducing_nodate}
Anthropic.
\newblock \href {https://www.anthropic.com/} {Introducing {Claude}}.

\bibitem[{Asch(2023)}]{asch_interview_2023}
David~A. Asch. 2023.
\newblock \href {https://doi.org/10.1056/CAT.23.0043} {An {Interview} with
  {ChatGPT} {About} {Health} {Care}}.
\newblock \emph{Catalyst non-issue content}, 4(2).
\newblock Publisher: Massachusetts Medical Society.

\bibitem[{Aydın and Karaarslan(2022)}]{aydin_openai_2022}
Ömer Aydın and Enis Karaarslan. 2022.
\newblock \href {https://doi.org/10.2139/ssrn.4308687} {{OpenAI} {ChatGPT}
  {Generated} {Literature} {Review}: {Digital} {Twin} in {Healthcare}}.

\bibitem[{Biswas(2023)}]{biswas_chatgpt_2023}
Som Biswas. 2023.
\newblock \href {https://doi.org/10.1148/radiol.223312} {{ChatGPT} and the
  {Future} of {Medical} {Writing}}.
\newblock \emph{Radiology}, 307(2):e223312.
\newblock Publisher: Radiological Society of North America.

\bibitem[{Cascella et~al.(2023)Cascella, Montomoli, Bellini, and
  Bignami}]{cascella_evaluating_2023}
Marco Cascella, Jonathan Montomoli, Valentina Bellini, and Elena Bignami. 2023.
\newblock \href {https://doi.org/10.1007/s10916-023-01925-4} {Evaluating the
  {Feasibility} of {ChatGPT} in {Healthcare}: {An} {Analysis} of {Multiple}
  {Clinical} and {Research} {Scenarios}}.
\newblock \emph{Journal of Medical Systems}, 47(1):33.

\bibitem[{Chen and Decary(2019)}]{chen_embedding_2019}
Mei Chen and Michel Decary. 2019.
\newblock \href {https://doi.org/10.3233/978-1-61499-951-5-59} {Embedding
  {Health} {Literacy} {Tools} in {Patient} {EHR} {Portals} to {Facilitate}
  {Productive} {Patient} {Engagement}}.
\newblock In \emph{Improving {Usability}, {Safety} and {Patient} {Outcomes}
  with {Health} {Information} {Technology}}, pages 59--63. IOS Press.

\bibitem[{Christiano et~al.(2023)Christiano, Leike, Brown, Martic, Legg, and
  Amodei}]{christiano_deep_2023}
Paul Christiano, Jan Leike, Tom~B. Brown, Miljan Martic, Shane Legg, and Dario
  Amodei. 2023.
\newblock \href {https://doi.org/10.48550/arXiv.1706.03741} {Deep reinforcement
  learning from human preferences}.
\newblock ArXiv:1706.03741 [cs, stat].

\bibitem[{Ely et~al.(2005)Ely, Osheroff, Chambliss, Ebell, and
  Rosenbaum}]{ely_answering_2005}
John~W. Ely, Jerome~A. Osheroff, M.~Lee Chambliss, Mark~H. Ebell, and Marcy~E.
  Rosenbaum. 2005.
\newblock \href {https://doi.org/10.1197/jamia.M1608} {Answering {Physicians}'
  {Clinical} {Questions}: {Obstacles} and {Potential} {Solutions}}.
\newblock \emph{Journal of the American Medical Informatics Association :
  JAMIA}, 12(2):217--224.

\bibitem[{Eysenbach(2023)}]{eysenbach_role_2023}
Gunther Eysenbach. 2023.
\newblock \href {https://doi.org/10.2196/46885} {The {Role} of {ChatGPT},
  {Generative} {Language} {Models}, and {Artificial} {Intelligence} in
  {Medical} {Education}: {A} {Conversation} {With} {ChatGPT} and a {Call} for
  {Papers}}.
\newblock \emph{JMIR Medical Education}, 9(1):e46885.
\newblock Company: JMIR Medical Education Distributor: JMIR Medical Education
  Institution: JMIR Medical Education Label: JMIR Medical Education Publisher:
  JMIR Publications Inc., Toronto, Canada.

\bibitem[{Fan et~al.(2023)Fan, Li, Ma, Lee, Yu, and
  Hemphill}]{fan_bibliometric_2023}
Lizhou Fan, Lingyao Li, Zihui Ma, Sanggyu Lee, Huizi Yu, and Libby Hemphill.
  2023.
\newblock \href {https://doi.org/10.48550/arXiv.2304.02020} {A {Bibliometric}
  {Review} of {Large} {Language} {Models} {Research} from 2017 to 2023}.
\newblock ArXiv:2304.02020 [cs].

\bibitem[{Gilson et~al.(2023)Gilson, Safranek, Huang, Socrates, Chi, Taylor,
  and Chartash}]{gilson_how_2023}
Aidan Gilson, Conrad~W. Safranek, Thomas Huang, Vimig Socrates, Ling Chi,
  Richard~Andrew Taylor, and David Chartash. 2023.
\newblock \href {https://doi.org/10.2196/45312} {How {Does} {ChatGPT} {Perform}
  on the {United} {States} {Medical} {Licensing} {Examination}? {The}
  {Implications} of {Large} {Language} {Models} for {Medical} {Education} and
  {Knowledge} {Assessment}}.
\newblock \emph{JMIR medical education}, 9:e45312.

\bibitem[{Hanauer et~al.(2015)Hanauer, Mei, Law, Khanna, and
  Zheng}]{hanauer_supporting_2015}
David~A. Hanauer, Qiaozhu Mei, James Law, Ritu Khanna, and Kai Zheng. 2015.
\newblock \href {https://doi.org/10.1016/j.jbi.2015.05.003} {Supporting
  information retrieval from electronic health records: {A} report of
  {University} of {Michigan}'s nine-year experience in developing and using the
  {Electronic} {Medical} {Record} {Search} {Engine} ({EMERSE})}.
\newblock \emph{Journal of Biomedical Informatics}, 55:290--300.

\bibitem[{Hosseini et~al.(2023)Hosseini, Gao, Liebovitz, Carvalho, Ahmad, Luo,
  MacDonald, Holmes, and Kho}]{hosseini_exploratory_2023}
Mohammad Hosseini, Catherine~A. Gao, David Liebovitz, Alexandre Carvalho,
  Faraz~S. Ahmad, Yuan Luo, Ngan MacDonald, Kristi Holmes, and Abel Kho. 2023.
\newblock \href {https://doi.org/10.1101/2023.03.31.23287979} {An exploratory
  survey about using {ChatGPT} in education, healthcare, and research}.
\newblock Pages: 2023.03.31.23287979.

\bibitem[{Hu and Hu(2023)}]{hu_chatgpt_2023}
Krystal Hu and Krystal Hu. 2023.
\newblock \href
  {https://www.reuters.com/technology/chatgpt-sets-record-fastest-growing-user-base-analyst-note-2023-02-01/}
  {{ChatGPT} sets record for fastest-growing user base - analyst note}.
\newblock \emph{Reuters}.

\bibitem[{Jeblick et~al.(2022)Jeblick, Schachtner, Dexl, Mittermeier, Stüber,
  Topalis, Weber, Wesp, Sabel, Ricke, and Ingrisch}]{jeblick_chatgpt_2022}
Katharina Jeblick, Balthasar Schachtner, Jakob Dexl, Andreas Mittermeier,
  Anna~Theresa Stüber, Johanna Topalis, Tobias Weber, Philipp Wesp, Bastian
  Sabel, Jens Ricke, and Michael Ingrisch. 2022.
\newblock \href {https://doi.org/10.48550/arXiv.2212.14882} {{ChatGPT} {Makes}
  {Medicine} {Easy} to {Swallow}: {An} {Exploratory} {Case} {Study} on
  {Simplified} {Radiology} {Reports}}.
\newblock ArXiv:2212.14882 [cs].

\bibitem[{Khairat et~al.(2020)Khairat, Coleman, Ottmar, Jayachander, Bice, and
  Carson}]{khairat_association_2020}
Saif Khairat, Cameron Coleman, Paige Ottmar, Dipika~Irene Jayachander, Thomas
  Bice, and Shannon~S. Carson. 2020.
\newblock \href {https://doi.org/10.1001/jamanetworkopen.2020.7385}
  {Association of {Electronic} {Health} {Record} {Use} {With} {Physician}
  {Fatigue} and {Efficiency}}.
\newblock \emph{JAMA Network Open}, 3(6):e207385.

\bibitem[{Khan et~al.(2023)Khan, Jawaid, Khan, and Sajjad}]{khan_chatgpt_2023}
Rehan~Ahmed Khan, Masood Jawaid, Aymen~Rehan Khan, and Madiha Sajjad. 2023.
\newblock \href {https://doi.org/10.12669/pjms.39.2.7653} {{ChatGPT} -
  {Reshaping} medical education and clinical management}.
\newblock \emph{Pakistan Journal of Medical Sciences}, 39(2):605--607.

\bibitem[{Kruskal and Wallis(1952)}]{kruskal_use_1952}
William~H. Kruskal and W.~Allen Wallis. 1952.
\newblock \href {https://doi.org/10.2307/2280779} {Use of {Ranks} in
  {One}-{Criterion} {Variance} {Analysis}}.
\newblock \emph{Journal of the American Statistical Association},
  47(260):583--621.
\newblock Publisher: [American Statistical Association, Taylor \& Francis,
  Ltd.].

\bibitem[{Kung et~al.(2023)Kung, Cheatham, Medenilla, Sillos, Leon, Elepaño,
  Madriaga, Aggabao, Diaz-Candido, Maningo, and Tseng}]{kung_performance_2023}
Tiffany~H. Kung, Morgan Cheatham, Arielle Medenilla, Czarina Sillos, Lorie~De
  Leon, Camille Elepaño, Maria Madriaga, Rimel Aggabao, Giezel Diaz-Candido,
  James Maningo, and Victor Tseng. 2023.
\newblock \href {https://doi.org/10.1371/journal.pdig.0000198} {Performance of
  {ChatGPT} on {USMLE}: {Potential} for {AI}-assisted medical education using
  large language models}.
\newblock \emph{PLOS Digital Health}, 2(2):e0000198.
\newblock Publisher: Public Library of Science.

\bibitem[{Lee()}]{lee_rise_nodate}
Hyunsu Lee.
\newblock \href {https://doi.org/10.1002/ase.2270} {The rise of {ChatGPT}:
  {Exploring} its potential in medical education}.
\newblock \emph{Anatomical Sciences Education}, n/a(n/a).
\newblock \_eprint: https://onlinelibrary.wiley.com/doi/pdf/10.1002/ase.2270.

\bibitem[{Mutabazi et~al.(2021)Mutabazi, Ni, Tang, and
  Cao}]{mutabazi_review_2021}
Emmanuel Mutabazi, Jianjun Ni, Guangyi Tang, and Weidong Cao. 2021.
\newblock \href {https://doi.org/10.3390/app11125456} {A {Review} on {Medical}
  {Textual} {Question} {Answering} {Systems} {Based} on {Deep} {Learning}
  {Approaches}}.
\newblock \emph{Applied Sciences}, 11(12):5456.
\newblock Number: 12 Publisher: Multidisciplinary Digital Publishing Institute.

\bibitem[{Ouyang et~al.(2022)Ouyang, Wu, Jiang, Almeida, Wainwright, Mishkin,
  Zhang, Agarwal, Slama, Ray, Schulman, Hilton, Kelton, Miller, Simens, Askell,
  Welinder, Christiano, Leike, and Lowe}]{ouyang_training_2022}
Long Ouyang, Jeff Wu, Xu~Jiang, Diogo Almeida, Carroll~L. Wainwright, Pamela
  Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John
  Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda
  Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022.
\newblock \href {https://doi.org/10.48550/arXiv.2203.02155} {Training language
  models to follow instructions with human feedback}.
\newblock ArXiv:2203.02155 [cs].

\bibitem[{PhysioNet(2023)}]{physionet_responsible_2023}
PhysioNet. 2023.
\newblock \href {https://physionet.org/news/post/415} {Responsible use of
  {MIMIC} data with online services like {GPT}}.

\bibitem[{Rae et~al.(2022)Rae, Borgeaud, Cai, Millican, Hoffmann, Song,
  Aslanides, Henderson, Ring, Young, Rutherford, Hennigan, Menick, Cassirer,
  Powell, Driessche, Hendricks, Rauh, Huang, Glaese, Welbl, Dathathri, Huang,
  Uesato, Mellor, Higgins, Creswell, McAleese, Wu, Elsen, Jayakumar,
  Buchatskaya, Budden, Sutherland, Simonyan, Paganini, Sifre, Martens, Li,
  Kuncoro, Nematzadeh, Gribovskaya, Donato, Lazaridou, Mensch, Lespiau,
  Tsimpoukelli, Grigorev, Fritz, Sottiaux, Pajarskas, Pohlen, Gong, Toyama,
  d'Autume, Li, Terzi, Mikulik, Babuschkin, Clark, Casas, Guy, Jones, Bradbury,
  Johnson, Hechtman, Weidinger, Gabriel, Isaac, Lockhart, Osindero, Rimell,
  Dyer, Vinyals, Ayoub, Stanway, Bennett, Hassabis, Kavukcuoglu, and
  Irving}]{rae_scaling_2022}
Jack~W. Rae, Sebastian Borgeaud, Trevor Cai, Katie Millican, Jordan Hoffmann,
  Francis Song, John Aslanides, Sarah Henderson, Roman Ring, Susannah Young,
  Eliza Rutherford, Tom Hennigan, Jacob Menick, Albin Cassirer, Richard Powell,
  George van~den Driessche, Lisa~Anne Hendricks, Maribeth Rauh, Po-Sen Huang,
  Amelia Glaese, Johannes Welbl, Sumanth Dathathri, Saffron Huang, Jonathan
  Uesato, John Mellor, Irina Higgins, Antonia Creswell, Nat McAleese, Amy Wu,
  Erich Elsen, Siddhant Jayakumar, Elena Buchatskaya, David Budden, Esme
  Sutherland, Karen Simonyan, Michela Paganini, Laurent Sifre, Lena Martens,
  Xiang~Lorraine Li, Adhiguna Kuncoro, Aida Nematzadeh, Elena Gribovskaya,
  Domenic Donato, Angeliki Lazaridou, Arthur Mensch, Jean-Baptiste Lespiau,
  Maria Tsimpoukelli, Nikolai Grigorev, Doug Fritz, Thibault Sottiaux, Mantas
  Pajarskas, Toby Pohlen, Zhitao Gong, Daniel Toyama, Cyprien de~Masson
  d'Autume, Yujia Li, Tayfun Terzi, Vladimir Mikulik, Igor Babuschkin, Aidan
  Clark, Diego de~Las Casas, Aurelia Guy, Chris Jones, James Bradbury, Matthew
  Johnson, Blake Hechtman, Laura Weidinger, Iason Gabriel, William Isaac,
  Ed~Lockhart, Simon Osindero, Laura Rimell, Chris Dyer, Oriol Vinyals, Kareem
  Ayoub, Jeff Stanway, Lorrayne Bennett, Demis Hassabis, Koray Kavukcuoglu, and
  Geoffrey Irving. 2022.
\newblock \href {https://doi.org/10.48550/arXiv.2112.11446} {Scaling {Language}
  {Models}: {Methods}, {Analysis} \& {Insights} from {Training} {Gopher}}.
\newblock ArXiv:2112.11446 [cs].

\bibitem[{Raghavan et~al.(2018)Raghavan, Patwardhan, Liang, and
  Devarakonda}]{raghavan_annotating_2018}
Preethi Raghavan, Siddharth Patwardhan, Jennifer~J. Liang, and Murthy~V.
  Devarakonda. 2018.
\newblock \href {https://doi.org/10.48550/arXiv.1805.06816} {Annotating
  {Electronic} {Medical} {Records} for {Question} {Answering}}.
\newblock ArXiv:1805.06816 [cs].

\bibitem[{Reshmi and Balakrishnan(2016)}]{reshmi_implementation_2016}
S.~Reshmi and Kannan Balakrishnan. 2016.
\newblock \href {https://doi.org/10.1007/s12046-016-0544-1} {Implementation of
  an inquisitive chatbot for database supported knowledge bases}.
\newblock \emph{Sādhanā}, 41(10):1173--1178.

\bibitem[{Roberts and Demner-Fushman(2015)}]{roberts_toward_2015}
Kirk Roberts and Dina Demner-Fushman. 2015.
\newblock Toward a natural language interface for {EHR} questions.
\newblock \emph{AMIA Summits on Translational Science Proceedings}, 2015:157.
\newblock Publisher: American Medical Informatics Association.

\bibitem[{Roberts et~al.(2016)Roberts, Demner-Fushman, Voorhees, and
  Hersh}]{roberts_overview_2016}
Kirk Roberts, Dina Demner-Fushman, Ellen~M Voorhees, and William~R Hersh. 2016.
\newblock Overview of the {TREC} 2016 {Clinical} {Decision} {Support} {Track}.
\newblock In \emph{Text {REtrieval} {Conference} ({TREC})}.

\bibitem[{Roman et~al.(2017)Roman, Ancker, Johnson, and
  Senathirajah}]{roman_navigation_2017}
Lisette~C. Roman, Jessica~S. Ancker, Stephen~B. Johnson, and Yalini
  Senathirajah. 2017.
\newblock \href {https://doi.org/10.1016/j.jbi.2017.01.005} {Navigation in the
  electronic health record: {A} review of the safety and usability literature}.
\newblock \emph{Journal of Biomedical Informatics}, 67:69--79.

\bibitem[{Ruby(2023)}]{ruby_57_2023}
Daniel Ruby. 2023.
\newblock \href {https://www.demandsage.com/chatgpt-statistics/} {57+ {ChatGPT}
  {Statistics} for 2023 ({New} {Data} + {GPT}-4 {Facts})}.

\bibitem[{Sallam(2023{\natexlab{a}})}]{sallam_chatgpt_2023}
Malik Sallam. 2023{\natexlab{a}}.
\newblock \href {https://doi.org/10.3390/healthcare11060887} {{ChatGPT}
  {Utility} in {Healthcare} {Education}, {Research}, and {Practice}:
  {Systematic} {Review} on the {Promising} {Perspectives} and {Valid}
  {Concerns}}.
\newblock \emph{Healthcare}, 11(6):887.
\newblock Number: 6 Publisher: Multidisciplinary Digital Publishing Institute.

\bibitem[{Sallam(2023{\natexlab{b}})}]{sallam_utility_2023}
Malik Sallam. 2023{\natexlab{b}}.
\newblock \href {https://doi.org/10.1101/2023.02.19.23286155} {The {Utility} of
  {ChatGPT} as an {Example} of {Large} {Language} {Models} in {Healthcare}
  {Education}, {Research} and {Practice}: {Systematic} {Review} on the {Future}
  {Perspectives} and {Potential} {Limitations}}.
\newblock Pages: 2023.02.19.23286155.

\bibitem[{Shen et~al.(2023)Shen, Heacock, Elias, Hentel, Reig, Shih, and
  Moy}]{shen_chatgpt_2023}
Yiqiu Shen, Laura Heacock, Jonathan Elias, Keith~D. Hentel, Beatriu Reig,
  George Shih, and Linda Moy. 2023.
\newblock \href {https://doi.org/10.1148/radiol.230163} {{ChatGPT} and {Other}
  {Large} {Language} {Models} {Are} {Double}-edged {Swords}}.
\newblock \emph{Radiology}, 307(2):e230163.
\newblock Publisher: Radiological Society of North America.

\bibitem[{Shneiderman et~al.(2013)Shneiderman, Plaisant, and
  Hesse}]{shneiderman_improving_2013}
Ben Shneiderman, Catherine Plaisant, and Bradford~W. Hesse. 2013.
\newblock \href {https://doi.org/10.1109/MC.2013.38} {Improving {Healthcare}
  with {Interactive} {Visualization}}.
\newblock \emph{Computer}, 46(5):58--66.
\newblock Conference Name: Computer.

\bibitem[{Shoeybi et~al.(2020)Shoeybi, Patwary, Puri, LeGresley, Casper, and
  Catanzaro}]{shoeybi_megatron-lm_2020}
Mohammad Shoeybi, Mostofa Patwary, Raul Puri, Patrick LeGresley, Jared Casper,
  and Bryan Catanzaro. 2020.
\newblock \href {http://arxiv.org/abs/1909.08053} {Megatron-{LM}: {Training}
  {Multi}-{Billion} {Parameter} {Language} {Models} {Using} {Model}
  {Parallelism}}.
\newblock ArXiv:1909.08053 [cs].

\bibitem[{Siad(2023)}]{siad_promise_2023}
S.~M. Siad. 2023.
\newblock \href {https://doi.org/10.17613/yb4n-mc79} {The {Promise} and
  {Perils} of {Google}'s {Bard} for {Scientific} {Research}}.

\bibitem[{Soni et~al.(2023)Soni, Datta, and Roberts}]{soni_quehry_2023}
Sarvesh Soni, Surabhi Datta, and Kirk Roberts. 2023.
\newblock \href {https://doi.org/10.1093/jamia/ocad050} {{quEHRy}: a question
  answering system to query electronic health records}.
\newblock \emph{Journal of the American Medical Informatics Association},
  30(6):1091--1102.

\bibitem[{Tawfik et~al.(2023)Tawfik, Ghallab, and Moustafa}]{tawfik_nurse_2023}
Elham Tawfik, Eman Ghallab, and Amel Moustafa. 2023.
\newblock \href {https://doi.org/10.1186/s12912-023-01243-7} {A nurse versus a
  chatbot ‒ the effect of an empowerment program on chemotherapy-related side
  effects and the self-care behaviors of women living with breast {Cancer}: a
  randomized controlled trial}.
\newblock \emph{BMC Nursing}, 22(1):102.

\bibitem[{Vaishya et~al.(2023)Vaishya, Misra, and Vaish}]{vaishya_chatgpt_2023}
Raju Vaishya, Anoop Misra, and Abhishek Vaish. 2023.
\newblock \href {https://doi.org/10.1016/j.dsx.2023.102744} {{ChatGPT}: {Is}
  this version good for healthcare and research?}
\newblock \emph{Diabetes \& Metabolic Syndrome: Clinical Research \& Reviews},
  17(4):102744.

\bibitem[{Zhang and Walji(2014)}]{zhang_better_2014}
Jiajie Zhang and Muhammad Walji. 2014.
\newblock \href {https://doi.org/10.13140/2.1.1921.1841} {\emph{Better {EHR}:
  {Usability}, {Workflow}, and {Cognitive} {Support} in {Electronic} {Health}
  {Records}}}.

\bibitem[{Zhang et~al.(2022)Zhang, Roller, Goyal, Artetxe, Chen, Chen, Dewan,
  Diab, Li, Lin, Mihaylov, Ott, Shleifer, Shuster, Simig, Koura, Sridhar, Wang,
  and Zettlemoyer}]{zhang_opt_2022}
Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui
  Chen, Christopher Dewan, Mona Diab, Xian Li, Xi~Victoria Lin, Todor Mihaylov,
  Myle Ott, Sam Shleifer, Kurt Shuster, Daniel Simig, Punit~Singh Koura, Anjali
  Sridhar, Tianlu Wang, and Luke Zettlemoyer. 2022.
\newblock \href {http://arxiv.org/abs/2205.01068} {{OPT}: {Open} {Pre}-trained
  {Transformer} {Language} {Models}}.
\newblock ArXiv:2205.01068 [cs].

\bibitem[{Zhao et~al.(2023)Zhao, Zhou, Li, Tang, Wang, Hou, Min, Zhang, Zhang,
  Dong, Du, Yang, Chen, Chen, Jiang, Ren, Li, Tang, Liu, Liu, Nie, and
  Wen}]{zhao_survey_2023}
Wayne~Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou,
  Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, Yifan Du, Chen Yang,
  Yushuo Chen, Zhipeng Chen, Jinhao Jiang, Ruiyang Ren, Yifan Li, Xinyu Tang,
  Zikang Liu, Peiyu Liu, Jian-Yun Nie, and Ji-Rong Wen. 2023.
\newblock \href {https://doi.org/10.48550/arXiv.2303.18223} {A {Survey} of
  {Large} {Language} {Models}}.
\newblock ArXiv:2303.18223 [cs].

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{acl}
\usepackage{multirow}
\usepackage{pgfplots}
\pgfplotsset{compat=1.18}
\usepackage{setspace}
\usepackage{comment}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\title{Evaluation of AI Chatbots for Patient-Specific EHR Questions}

\author{Alaleh Hamidi, Kirk Roberts \\
  McWilliams School of Biomedical Informatics \\
  The University of Texas Health Science Center at Houston \\
  \texttt{\{alaleh.hamidi,kirk.roberts\}@uth.tmc.edu} \\}

\begin{document}
\maketitle
\begin{abstract}
This paper investigates the use of artificial intelligence chatbots for patient-specific question answering (QA) from clinical notes using several large language model (LLM) based systems: ChatGPT (versions 3.5 and 4), Google Bard, and Claude. We evaluate the accuracy, relevance, comprehensiveness, and coherence of the answers generated by each model using a 5-point Likert scale on a set of patient-specific questions.

At present time, we report results for ChatGPT 3.5 and Claude. The results indicate that both models are able to provide accurate, relevant, and comprehensive answers to a wide range of questions. 

Overall, our results suggest that LLMs are a promising tool for patient-specific QA from clinical notes. Further research is needed to improve the comprehensiveness and coherence of the answers generated by LLMs, and to investigate the use of LLMs in other medical applications.
\end{abstract}

\section{Introduction}

Automatic question answering (QA) systems, such as those based on large language models (LLMs), have made significant progress across a wide range of QA tasks, particularly in the medical field \cite{mutabazi_review_2021}.

LLMs are a class of language models that have shown exceptional performance across a variety of natural language processing (NLP) tasks due to their capacity to model and generate human-like language \cite{fan_bibliometric_2023}. LLMs utilize neural networks with billions of parameters and are trained using self-supervised learning with substantial volumes of unlabeled text data \cite{shen_chatgpt_2023, zhao_survey_2023}. GPT-3 and 4, Google Bard \cite{siad_promise_2023}, Gopher \cite{rae_scaling_2022}, Megatron \cite{shoeybi_megatron-lm_2020}, and OPT-175B \cite{zhang_opt_2022} are only a few examples of LLMs.

ChatGPT is a successor of InstructGPT \cite{ouyang_training_2022} with a fine-tuned dialog interface using Reinforcement Learning with Human Feedback (RLHF) \cite{christiano_deep_2023}. It is trained on approximately 100 trillion parameters and 300 billion words (only till September 2021) and launched on November 30, 2022, as a tool to allow users to converse with a machine about various subjects. Having crossed 1 billion users in March 2023 \cite{ruby_57_2023}, ChatGPT has set records for consumer interest in artificial intelligence (AI) systems \cite{hu_chatgpt_2023}.

Bard is another LLM-based chatbot developed by Google \cite{siad_promise_2023} using the LaMDA (Language Model for Dialogue Applications) architecture and released in March 2023. Bard is also trained on massive amounts of data, including books, articles, and other written text, to generate text, translate languages, write different kinds of creative content, and answer questions. It is designed to deliver real-time information from a wide range of sources on the Internet. Bard is still under development, but it has the potential to be a powerful tool for a variety of applications.  \cite{ noauthor_google_2023}.

Claude is comparatively smaller than many other models, having less than 100 million parameters, compared to LLMs like OpenAI's GPT-3, with over 10 billion parameters. Claude was launched in March 2023 by Anthropic to ensure safety and ethics. Based on Anthropics's claim, this chatbot's focus is on being helpful, harmless, and honest using a technique called Constitutional AI which has not been publicly disclosed. In other words, Claude does not do unconstrained, open-domain text generation, which is a key feature of most LLMs \cite{anthropic_introducing_nodate}. 

The potential application of language models in health education, research, and practice can be promising, albeit if the legitimate issues are proactively considered and addressed \cite{sallam_chatgpt_2023}. In terms of their medical applications, patient-specific QA from clinical notes is an essential but challenging task that can provide clinicians with quicker access to patient information to facilitate medical decision-making \cite{soni_quehry_2023, roberts_toward_2015, raghavan_annotating_2018}. However, most existing work on medical QA applications has concentrated on evaluating and extracting general medical knowledge, such as medical licensing exams \cite{kung_performance_2023, ali_performance_2023}, interviews \cite{asch_interview_2023}, and education \cite{ eysenbach_role_2023, sallam_utility_2023, hosseini_exploratory_2023, khan_chatgpt_2023, lee_rise_nodate}.
Meanwhile, using the language models for patient-specific QA from electronic health records (EHRs) has gained comparatively little study. Although there are some works in applying recent LLMs in patient-specific knowledge from EHRs \cite{jeblick_chatgpt_2022}, they are not question answering.

This work studies how accurate and comprehensive several popular AI chatbots are at answering medical questions based on patients' clinical notes in two different prompt scenarios. Based on a publicly-available subset of MIMIC-III notes released as part of the TREC 2016 Clinical Decision Support track\footnote{Note that in general MIMIC notes are access-restricted \cite{physionet_responsible_2023}. We use a special subset for which public release was granted in 2016.} \cite{roberts_overview_2016}, which includes 30 de-identified clinical notes, we consider ChatGPT (versions 3.5 and 4), Google Bard, and Claude to understand how accurate, relevant, comprehensive, and coherent the answers are for each language model.

\section{Related Work}

Clinicians frequently access electronic health records (EHRs) to obtain valuable patient information for providing care. However, the usability \cite{zhang_better_2014} and navigation \cite{roman_navigation_2017} challenges associated with EHRs impede the efficient retrieval of information from these systems \cite{khairat_association_2020}. Current approaches to address these challenges involve visualizing information as charts \cite{shneiderman_improving_2013} or using keyword-based searches for information retrieval (IR) \cite{hanauer_supporting_2015}. While these methods enhance information access, they still overwhelm users with excessive details, such as presenting a table of lab values or a lengthy list of procedures, when only a specific lab value or status is required. In other words, these methods fail to understand the precise information needs of users accurately. Consequently, question answering (QA) emerges as a natural solution, offering a means to identify information needs and provide an exact, verifiable answer \cite{ely_answering_2005}.

A QA system can interpret and comprehend questions posed by users using everyday language. It possesses a comprehensive information retrieval capability to provide answers. Typically, QA systems consist of three main components: question analysis, information retrieval, and answer generation. Chatbots designed to simulate human conversation, allowing interactions and dialogue using natural language including substantial reliance on QA functionality \cite{reshmi_implementation_2016}. AI chatbots have been employed to provide health advice. Studies have shown that AI chatbots may enhance self-care and lead to improved patient outcomes \cite{tawfik_nurse_2023, chen_embedding_2019}. ChatGPT and Google Bard are AI chatbots that have been researched substantially for their feasibility and application in the medical domain. The feasibility of ChatGPT has been investigated in clinical and research scenarios such as support of the clinical practice, scientific production (such as creating a literature review article \cite{aydin_openai_2022} or writing a scientific paper \cite{biswas_chatgpt_2023}), misuse in medicine and research, and reasoning about public health topics \cite{cascella_evaluating_2023, sallam_utility_2023, vaishya_chatgpt_2023}.

Although some studies employ AI chatbots to process EHRs and evaluate the results (e.g., assessing the quality of radiology reports simplified by ChatGPT \cite{jeblick_chatgpt_2022}), they do not use the chatbot as an interactive QA tool to obtain specific information regarding the patient. On the other hand, the works which utilize AI chatbots as a conversational tool do not consider patient-specific notes. For instance, \newcite{kung_performance_2023} and \newcite{gilson_how_2023} both evaluated the performance of ChatGPT on USMLE, the United States Medical Licensing Exam \cite{kung_performance_2023}.


The lack of research in the evaluation of AI chatbots for medical QA based on patient-specific notes is the key motivation for conducting this study.

\section{Methodology}
In this work, we use admission notes from MIMIC-III (taken from the TREC CDS 2016 topics \cite{roberts_overview_2016}) to evaluate ChatGPT (3.5 \& 4), Google Bard, and Claude. Admission notes are created by clinicians, primarily physicians, to describe a patient's principal complaint, pertinent medical history, and any other details obtained during the initial hours of their hospital stay, such as lab results. Substantial numbers of abbreviations as well as other linguistic jargon and style are used in these notes. The notes used in this study have all been de-identified to protect patient privacy. The first step in this work was therefore substituting the de-identification labels with appropriate values to have human-like notes.

For each clinical note, three categories of questions with five questions in each category were created:
\begin{itemize}
    \item \textbf{General Questions}: Could apply to almost any patient (age, sex, reason for admission, etc.) and answerable from a specific text span.
    \item \textbf{Specific Questions}: Targeted to the information in the note itself and answerable from a specific text span.
    \item \textbf{Nonanswerable Questions}: Targeted to the information in the note, but not actually answerable from the note itself.
\end{itemize}

Since chatbots are dialogue systems and maintain state over the course of a session, two different scenarios were considered:

\begin{itemize}
    \item \textbf{1 Question per Session (1QpS)}: In this scenario, the chatbot is presented with the patient's note and one question per session. No state about prior questions for the patient note is kept.
    
    \item \textbf{1 Topic per Session (1TpS)}: In this scenario, the chatbot is presented with the patient's note and all 15 questions are asked within the same session. In other words, the note and the history of previously asked questions with their given answers are kept. The questions are asked in a random order.
\end{itemize}

Then, individuals with medical backgrounds evaluated the answers in terms of accuracy, relevance, coverage, and coherence and assigned each answer a score between 1 (very weak) to 5 (excellent) for each evaluation criterion:

\begin{itemize}
    \item \textbf{Accuracy (Correctness)}: Is the information provided in the answer factually correct according to the note? Are there any incorrect or contradictory statements?
    \item \textbf{Relevance}: How relevant is the answer to the question? Does it address the key concepts and details asked in the question?
    \item \textbf{Coverage}: How comprehensive is the answer? Does it cover all relevant aspects, or is it lacking in some details?
    \item \textbf{Coherence}: Is the answer coherent and logically structured?

\end{itemize}

\section{Results}
At the present time we report just the evaluation of ChatGPT 3.5 and Claude in answering clinical questions from admission notes.
Additional experiments with other systems and more raters are ongoing.
Tables \ref{tab:Accuracy}- \ref{tab:Coherence} compare the performance of ChatGPT 3.5 and Claude in terms of accuracy, relevance, coverage, and coherence for different types of questions (General, Specific, Non-answerable) in two different scenarios (1 Question per Session and 1 Topic per Session). 

To study whether session scenarios, question type, or chatbot model have an effect on the answers' quality, Kruskal-Wallis analysis of variance \cite{kruskal_use_1952}, a non-parametric statistical test, was conducted.

Based on the results of the Kruskal-Wallis test, it cannot be concluded that there is a statistically significant difference in the average values in terms of the cases session scenario, question type, or AI model. The p-value for all these cases is greater than 0.05, and the null hypothesis (there is no statistically significant difference between the groups) cannot be rejected. In other words, the results indicate that both ChatGPT 3.5 and Claude performed well in answering questions based on admission notes. They provided accurate, relevant, comprehensive, and coherent answers across different question types and scenarios.

\begin{table*}
\centering
\begin{tabular}{| c  c  c  c  c c |}
\hline
\textbf{} & \textbf{} & \textbf{General } & \textbf{Specific} &\textbf{Non-answerable} &\textbf{Total}\\
\textbf{} & \textbf{} & \textbf{Questions} & \textbf{Questions} &\textbf{Questions} &\textbf{}\\
\hline
\hline
\multirow{2}{*}{\textbf{ChatGPT 3.5}}& \textbf{1QpS}  & 4.49 & 4.45 & 4.77 & 4.57\\
& \textbf{1TpS}  & 4.18 & 4.75 & 4.81 & 4.58\\
\hline
\multirow{2}{*}{\textbf{Claude}} & \textbf{1QpS}  & 4.52 & 4.66 & 4.23 & 4.47\\
& \textbf{1TpS}   & 4.61 & 4.60 & 4.58 & 4.60\\
\hline
\end{tabular}
\caption{Accuracy comparison of answers to three types of questions given to ChatGPT 3.5 and Claude in two different scenarios; one Question per Session and one Topic per Session}
\label{tab:Accuracy}
\end{table*}

\begin{table*}
\centering
\begin{tabular}{| c  c  c  c  c c |}
\hline
\textbf{} & \textbf{} & \textbf{General } & \textbf{Specific} &\textbf{Non-answerable} &\textbf{Total}\\
\textbf{} & \textbf{} & \textbf{Questions} & \textbf{Questions} &\textbf{Questions} &\textbf{}\\
\hline
\hline
\multirow{2}{*}{\textbf{ChatGPT 3.5}}& \textbf{1QpS}  & 4.77 & 4.75 & 4.86 & 4.79\\
& \textbf{1TpS}  & 4.49 & 4.91 & 4.83 & 4.74\\
\hline
\multirow{2}{*}{\textbf{Claude}} & \textbf{1QpS}  & 4.68 & 4.66 & 4.30 & 4.55\\
& \textbf{1TpS}   & 4.78 & 4.80 & 4.72 & 4.77\\
\hline
\end{tabular}
\caption{Relevance comparison of answers to three types of questions given to ChatGPT 3.5 and Claude in two different scenarios; one Question per Session and one Topic per Session}
\label{tab:Relevance}
\end{table*}

\begin{table*}
\centering
\begin{tabular}{| c  c  c  c  c c |}
\hline
\textbf{} & \textbf{} & \textbf{General } & \textbf{Specific} &\textbf{Non-answerable} &\textbf{Total}\\
\textbf{} & \textbf{} & \textbf{Questions} & \textbf{Questions} &\textbf{Questions} &\textbf{}\\
\hline
\hline
\multirow{2}{*}{\textbf{ChatGPT 3.5}}& \textbf{1QpS}  & 4.43 & 4.43 & 4.81 & 4.56\\
& \textbf{1TpS}  & 4.15 & 4.86 & 4.85 & 4.62\\
\hline
\multirow{2}{*}{\textbf{Claude}} & \textbf{1QpS}  & 4.80 & 4.78 & 4.75 & 4.78\\
& \textbf{1TpS}   & 4.72 & 4.66 & 4.77 & 4.72\\
\hline
\end{tabular}
\caption{Coverage comparison of answers to three types of questions given to ChatGPT 3.5 and Claude in two different scenarios; one Question per Session and one Topic per Session}
\label{tab:Coverage}
\end{table*}

\begin{table*}
\centering
\begin{tabular}{| c  c  c  c  c c |}
\hline
\textbf{} & \textbf{} & \textbf{General } & \textbf{Specific} &\textbf{Non-answerable} &\textbf{Total}\\
\textbf{} & \textbf{} & \textbf{Questions} & \textbf{Questions} &\textbf{Questions} &\textbf{}\\
\hline
\hline
\multirow{2}{*}{\textbf{ChatGPT 3.5}}& \textbf{1QpS}  & 4.58 & 4.83 & 4.86 & 4.76\\
& \textbf{1TpS}  & 4.32 & 4.77 & 4.81 & 4.63\\
\hline
\multirow{2}{*}{\textbf{Claude}} & \textbf{1QpS}  & 4.34 & 3.90 & 3.60 & 3.95\\
& \textbf{1TpS}   & 4.45 & 4.60 & 4.74 & 4.59\\
\hline
\end{tabular}
\caption{Coherence comparison of answers to three types of questions given to ChatGPT 3.5 and Claude in two different scenarios; one Question per Session and one Topic per Session}
\label{tab:Coherence}
\end{table*}

\section{Discussion}
The results of this study suggest that large language models (LLMs) can be used to generate accurate, relevant, comprehensive, and coherent answers to clinical questions from admission notes. This is a promising finding, as it suggests that LLMs could be used to provide clinicians with additional support in their work.
There are a few limitations to this study that should be noted. First, the data set used in this study was relatively small. It is possible that the results would be different if a larger data set were used. 
Second, the evaluation above has been conducted by one individual which carries a higher bias risk (evaluations with additional raters are ongoing).
Despite these limitations, the results of this study suggest that LLMs have the potential to be a valuable tool for clinicians. Further research is needed to evaluate the performance of LLMs on larger data sets and with a wider range of LLMs.

\section{Conclusion}
This study evaluates the performance of ChatGPT 3.5 and Claude in answering clinical questions based on MIMIC-III clinical notes. Both models show promising performance in the medical domain and exhibit substantial accuracy, relevance, coverage, and coherence in their responses. While further research and improvements are necessary to address certain limitations, these models hold great potential as valuable tools in the healthcare industry, assisting healthcare professionals in accessing relevant information and improving patient care. Future studies can identify the factors contributing to LLMs' performance in answering clinical questions.

\bibliography{LLMmodels_biblio}
\bibliographystyle{acl_natbib}

%
%

\end{document}